OpenStudy (anonymous):

find the 12th term of 1/2, 3/4, 9/8, ...

OpenStudy (akashdeepdeb):

Common ratio or r : 3/2 n=12 [Because we are finding the 12th term. First term or a = 1/2 nth term = ar^(n-1) Substitutethe values given in the expression and get the 12th term. Understood. :)

OpenStudy (akashdeepdeb):

r=3/2 okay?

OpenStudy (anonymous):

43.25 ?

OpenStudy (anonymous):

how did you get the common ratio r as 3/2

OpenStudy (anonymous):

that was my problem

OpenStudy (yttrium):

You can find the common ratio by dividing the 2nd term to the 1st term. For example, divide 3/4 by 1/2 This will be 3/4 * 2 Therefore, r = 3/2

OpenStudy (anonymous):

Is always in all cases?

OpenStudy (yttrium):

Yes. :)